Show that 224-1 and 216-1 are composite. Hints:use expansion of (a2-b2)


1
Expert's answer
2022-02-16T11:10:20-0500

(1) "\left(2^{24}-1\right)"

Now "a^{2}-b^{2}=(a-b)(a+b)"

"\begin{array}{rl}

\therefore 2^{24}-1=\left(2^{12}\right)^{2}-1^{2}=\left(2^{12}-1\right)\left(2^{12}+1\right) \\

=\left(\left(2^{6}\right)^{2}-1^{2}\right)\left(2^{12}+1\right) \\

=\left(2^{6}-1\right)\left(2^{6}+1\right)\left(2^{12}+1\right) \\

=\left(2^{3}-1\right)\left(2^{3}+1\right)\left(2^{6}+1\right)\left(2^{12}+1\right) \\

= 7 \times 9 \times\left(2^{6}+1\right)\left(2^{12}+1\right)

\end{array}"

"\therefore 2^{24}-1" has the divisor "7,9,2^{6}+1,2^{12}+1" other than the "1 \& 2^{24}-1" .

"\therefore 2^{24}-1" is composite number.

(2) "2^{16}-1"

"\begin{aligned}

2^{16}&-1=\left(2^{8}\right)^{2}-1^{2}=\left(2^{8}-1\right)\left(2^{8}+1\right) \\

&=\left(2^{4}-1\right)\left(2^{4}+1\right)\left(2^{8}+1\right) \\

&=\left(2^{2}-1\right)\left(2^{2}+1\right)\left(2^{4}+1\right)\left(2^{8}+1\right) \\

&=3 \times 5 \times 17 \times\left(2^{8}+1\right)

\end{aligned}"

 

"\therefore 2^{16}-1" has the divisor other than 1 & "2^{16}-1" .

"\therefore 2^{16}-1" is composite number.